Fermi National Laboratory


Accelerator Update - June 11-June 13

The integrated luminosity for last week was 4674.7 nb-1.

Wednesday June 11
The day shift began with the Antiproton Source (Pbar) stacking, with MiniBooNE taking beam, and with experts conducting Tevatron (TeV) and Switchyard studies.

At 12:33 PM, Operations halted stacking to cool the core before putting the TeV into shot setup.

Experts changed the controls for a Pbar Debuncher RF station (DRF1) at 1:18 PM as a test. Controls were then returned to normal.

During the morning a Linac RF station (KRF1) had been causing problems. At 2:58 PM, after Operations loaded protons into the TeV for the shot, Linac technicians began repairing the station.

At 3:26 PM, Operations established store 2673 with an initial luminosity of 31.47E30.

Booster vacuum valves closed at 4:26 PM, stopping all downstream beam. Operators had the valves open and beam restored five minutes later.

Booster next suffered trips on four RF stations (BRF3, 4, 5, & 6). They all reset okay.

At 10 PM, the Linac and Booster beam intensity had dropped. Operators tuned the beam.

Thursday June 12
The midnight shift began with Operations monitoring store 2673, with Pbar stacking, and with MiniBooNE taking beam.

Booster had two RF stations (BRF11 & 12) trip off at 2:24 AM. Operators reset BRF11, but didn't get BRF12 back on until 2:44 AM.

At 3:16 AM, a TeV RF station (TRF5) tripped off.

Problems with the Booster beam appeared again at 5:37 AM. Operators resumed tuning the beam, which of course affected the MiniBooNE beam. Then, at 6:39 AM, two RF stations (BRF13, & 14) tripped off. Operators reset the stations.

TeV experts began end-of-store studies at 7:31 AM.

Operations terminated store 2673 at 8:21 AM. TeV experts began a beam maintenance period and Switchyard experts began a study period.

Operations turned TRF5 back on at 8:46 AM.

At noon, Operations put the TeV into shot setup.

Booster suffered a vacuum burst at 1:08 PM. A vacuum tech began changing a vacuum pump power supply at 2:43 PM. Stacking and MiniBooNE beam resumed a half hour later.

At 3:18 PM, Operations established store 2676 with an initial luminosity of 27.54E30. Pbar resumed stacking.

Problems with Linac quadrupole controls and power supplies for QPS308 and 307 began at 6:22 PM. They continued until 7:33 PM when operators called in an expert to change out a power supply. The expert had QPS307 operational by 8:16 PM.

Booster beam problems reappeared at 10:27 PM. Operators resumed tuning.

The TeV suffered a sector D0 quench at 10:22 PM. Operations discovered a HFU had failed, which caused a heater to discharge and warm the magnet string.

Another Booster vacuum burst occurred at 11 PM.

At 11:50 PM, operators made the Booster ready for an access to look for a possible LCW leak. A Booster RF station (BRF4) had tripped off on a leak indication.

Friday June 13
The midnight shift began with operators accessing Booster to look for a LCW leak on BRF4, with a TeV cryo system recovering, and with operations monitoring the Pbar stack.

Operators returned from their Booster access at 12:56 AM. BRF4 simply needed a new aspirin; humidity had caused it to disintegrate. Aspirin tablets are used in spring style switches because they readily dissolve when hit with water.

At 1:33 AM, Operations resumed sending beam to Booster and MiniBooNE, but the Booster beam was very poor.

A TeV expert began a study period at 4:32 AM.

A Pbar Debuncher quadrupole power supply fault halted stacking at 6:56 AM. Operators managed to get the power supply working fifteen minutes later.

At 7:10 AM, the Pbar stack was lost. An Accumulator device (A:LQ) tripped off. Operators called in an expert.

Today's Plan
The plan for today is to repair A:LQ and resume stacking. During this downtime TeV experts will conduct beam maintenance. This afternoon the Collider Detector Facility and D-zero will be allowed a one-hour access.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/16/2003   email Fermilab